In a house four 20W lamps., two 10W and three 40W lamps are used 5hrs daily. Calculate the total power of the house loads and the energy consumed weekly?
1
Expert's answer
2020-04-23T12:45:00-0400

First of all lets's calculate the total power of the house

1) 20 W lamps x 4 =80 W

2) 10 W lamps x 2 =20 W

3) 40 W lamps x 3 =120 W

total power = 80 + 20 + 120 = 220 W

If each of the appliances are used 5 hours daily then in one week they are used 5*7=35 hours

energy consumed weekly = 220 x 35 = 7700 Wh = 7.7 kWh